a curve is such that d²y/dx²=12x-12. find the equation at point (2,9)?


1
Expert's answer
2022-02-25T12:37:34-0500
"\dfrac{dy}{dx}=\int(12x-12)dx=6x^2-12x+C_1"

"y=\int(6x^2-12x+C_1)dx=2x^3-6x^2+C_1 x+C_2"

Point "(2,9)"


"9=2(2)^3-6(2)^2+2C_1+C_2"

"C_2=17-2C_1"

The equation of the curve is


"y=2x^3-6x^2+C x+17-2C"